1.) 4b+2(b-4)=8(b+3)
2.) 8r+2=6r+10
3.) 1/5x-1/10x=7/2
4.) 8y-3-9y=5+y-12

Answer by rfer(16322)

4b+2b-8=8b+24
6b-8=8b+24
-2b=32
b=-16